Exhibit 16-4
In a laboratory experiment, data were gathered on the life span (Y in months) of 33 rats, units of daily protein intake (X1), and whether or not agent X2 (a proposed life extending agent) was added to the rats diet (X2 = 0 if agent X2 was not added, and X2 = 1 if agent was added.) From the results of the experiment, the following regression model was developed.



Also provided are SSR = 60 and SST = 180.

Refer to Exhibit 16-4. The degrees of freedom associated with SSE are

Group of answer choices

3

33

32

30
Answer #1

Total number of observation are 33

Total number of independent variable is 2

Total degree of freedom is n-1=33-1=32

Degree of freedom of error = 32-2=30

Option d is true